OpenStudy (anonymous):

multiply and simplify the rational expression to its lowest form. x^2/x^2+8 times x+1/x The answer I got is (x)(x+1)

OpenStudy (anonymous):

\[\frac{ x ^{2} }{ x ^{2}+8}.\frac{ x+1 }{ x}\]

OpenStudy (lgbasallote):

where did x^2 + 8 in your answer go?

OpenStudy (anonymous):

I figured you would break it down to (x+4)(x+4) ans they cancel themselves out

OpenStudy (lgbasallote):

\[\huge \implies \frac{x^\cancel 2}{x^2 + 8} \times \frac{x+1}{\cancel x} \implies \frac{x(x+1)}{x^2 + 8}\]

OpenStudy (lgbasallote):

and no.. x^2 + 8 is not equal to (x+4)(x+4)

OpenStudy (lgbasallote):

\[(x+4)(x+4) = x^2 + 4x + 4x + 16 \implies x^2 + 8x + 16\]

OpenStudy (anonymous):

Thank youfor the break down

OpenStudy (lgbasallote):

welcome